### v1

# Anisotropic capillary hypersurfaces in a wedge

Abstract.
We establish Minkowski formulae and a Heintze-Karcher type inequality for anisotropic hypersurfaces in a classical wedge in Euclidean space. As an application, we prove an Alexandrov-type theorem. This extends the results in [12] to the anisotropic case.
Key words and phrases:
Alexandrov Theorem, Heintze-Karcher inequality, constant mean curvature, capillary hypersurface, Wulff shape, wedge
2020 Mathematics Subject Classification:
53C24, 53C42, 53C40
1. Introduction
Alexandrov’s Soap Bubble Theorem [1] is a fundamental result in differential geometry, stating that any closed embedded hypersurface in $\mathbb{R}^{n+1}$ with constant mean curvature must be a round sphere. It establishes important links between the mean curvature and the rigidity of a closed hypersurface. The various methods to prove the Alexandrov theorem bring distinct perspectives and significant tools. For instance, Alexandrov’s approach involves reflecting through moving planes based on the maximum principle; Reilly [24] presented a new proof by his famous integration formula; Montiel and Ros [20] combined the Minkowski formula and the Heintze-Karcher inequality to offer a geometric proof; The proof by Hajazi, Montiel, and Zhang [11] uses a spinorial Reilly-type formula. Brendle [2] achieved a significant generalized Alexandrov theorem in warped product manifolds by employing the normal geodesic flow with respect to a conformal metric.
Capillary hypersurfaces refer to a fascinating area of study within the field of differential geometry that examines the shape and behavior of surfaces in the presence of capillary forces ([6, 7]). These forces emerge from the interaction between a liquid and a solid. The solid surfaces can take various shapes, such as wedges, cones or balls. Recently, there have been many new results in this field, including stability problems, overdetermined problems and the regularity of minimizing capillary hypersurfaces (refer to [3, 5, 8, 15, 17, 19, 25, 27] and the references therein).
About the Alexandrov theorem on capillary hypersurfaces, Wente initially studied the case of half-space +$\mathbb{R}^{n+1}_{+}$. Park [21] later obtained a similar result for the ring-type spanner in a wedge by Alexandrov’s reflection argument. Subsequently, Choe and Park [4] and López [18] considered the theorem in convex cone and wedge by the method of Reilly formula, respectively. Pyo [23] got rigidity theorems of hypersurfaces with free boundary in a wedge in a space form. Jia, Wang, Xia and Zhang [12] further obtained the Alexandrov theorem for capillary hypersurfaces in a wedge. More recently, Wang and Xia [26] introduced a novel perspective on capillary hypersurfaces in a unit ball by reducing them to free boundary cases via a special Finsler metric, thereby leading to the Alexandrov theorem.
For the anisotropic setting, He, Li, Ma and Ge [10] first proved the anisotropic version of the Alexandrov theorem for closed hypersurfaces in Euclidean space. Recently, Jia, Wang, Xia and Zhang [14] extended Wente’s result to anisotropic capillary hypersurfaces in the half-space +$\mathbb{R}^{n+1}_{+}$.
In this paper, we focus on anisotropic hypersurfaces in a classical wedge. Let $\bm{n}_{1},\bm{n}_{2}$ be two linearly independent unit vectors in $\mathbb{R}^{n+1}$. A wedge $\mathbf{W}$ determined by $\bm{n}_{1},\bm{n}_{2}$ is defined to be the set
$\{x\in\mathbb{R}^{n+1}\mid\langle x,\bm{n}_{i}\rangle<0,\,i=1,2\}.$
This is an open region in $\mathbb{R}^{n+1}$ bounded by two half-hyperplanes. The closure $\overline{\mathbf{W}}$ is a smooth manifold with corners. The boundary $\partial\mathbf{W}$ consists of two open half-hyperplanes $P_{1},P_{2}$ and a codimension $2$ linear subspace $L$ in $\mathbb{R}^{n+1}$. Here $L$ is the boundary of $P_{1}$ and $P_{2}$.
Now let $\Sigma\subset\overline{\mathbf{W}}$ be a smooth compact hypersurface in a classical wedge with $\partial\Sigma\subset\partial\mathbf{W}$ and $\Omega$ be the enclosed domain by $\Sigma$ and $\partial\mathbf{W}$. Following [14, 16], we define the energy functional:
$E(\Sigma)=\int_{\Sigma}F(\bm{\nu})\,\mathrm{d}A+\omega_{0}^{i}\left|\partial\Omega\cap P_{i}\right|$
where the first term $\int_{\Sigma}F(\bm{\nu})\,\mathrm{d}A$ represents the anisotropic surface tension. The second term $\omega_{0}^{i}\left|\partial\Omega\cap P_{i}\right|$ represents the wetting energy, where $i$ ranges from $1$ to $2$, and $\omega_{0}^{i}\in\mathbb{R}$ is a given constant.
Under a volume constraint, we consider the first variation of the energy functional $E$. For a family of hypersurfaces $\left\{\Sigma_{t}\right\}$ that vary smoothly, with boundaries $\partial\Sigma_{t}$ moving freely on the boundary $\partial\mathbf{W}$, and according to a variational vector field $Y$ such that $\left.Y\right|_{\partial\Sigma}\in T\left(\partial\mathbf{W}\right)$, the first variation formula of $E$ is given by
$\left.\frac{\mathrm{d}}{\mathrm{d}t}\right|_{t=0}E\left(\Sigma_{t}\right)=\int_{\Sigma}H^{F}\langle Y,\bm{\nu}\rangle\mathrm{d}A+\int_{\partial\Sigma\cap P_{i}}\langle Y,R_{i}(p_{i}(\Phi(\bm{\nu})))\rangle\mathrm{d}s-\omega_{0}^{i}\int_{\partial\Sigma\cap P_{i}}\langle Y,\mathbf{m}_{i}\rangle\mathrm{d}s,$
Here, $H^{F}$ denotes the anisotropic mean curvature of $\Sigma$, $p_{i}$ is the projection onto the $\left\{\bm{\nu},\bm{n}_{i}\right\}$-plane, $R_{i}$ is the $\pi/2$-rotation in the $\left\{\bm{\nu},\bm{n}_{i}\right\}$-plane and $\mathbf{m}_{i}$ is the conormal of $\partial\Sigma\subset P_{i}$. Furthermore,
$\Phi\colon\mathbb{S}^{n}\rightarrow\mathbb{R}^{n+1},\quad\Phi(x):=\nabla F(x)+F(x)x.$
By direct calculations, the first variational formula leads to
$H^{F}=const\,\text{ on }\Sigma\quad\text{ and }\quad\left\langle\Phi(\bm{\nu}),\bm{n}_{i}\right\rangle=\omega_{0}^{i}\quad\text{ on }\partial\Sigma\cap P_{i}.$
Inspired by this formula, we can define an anisotropic $\bm{\omega}_{0}$-capillary hypersurface in the classical wedge $\mathbf{W}$ as one that satisfies the following condition on its boundary:
$\left\langle\bm{\nu}^{F},\bm{n}_{i}\right\rangle=\omega^{i}_{0}\quad\text{ on }\partial\Sigma\cap P_{i},$
where $\bm{\omega}_{0}:=(\omega^{1}_{0},\omega^{2}_{0})$ is a constant vector and $\bm{\nu}^{F}:=\Phi(\bm{\nu})$ is the anisotropic normal vector field of $\Sigma$.
Our motivation in this paper is to present an Alexandrov type theorem. Building upon the geometric proof by Montiel and Ros [20], we aim to integrate the Minkowski formulae and the Heintze-Karcher inequality. As is well known, the Minkowski formula for closed hypersurfaces in Euclidean space can be derived by applying the divergence theorem to the tangential component of the position vector field. It is very interesting to observe that by applying the divergence theorem to the vector field below, we can derive the Minkowski formulae in a wedge, where
$X(x)=\langle\bm{\nu}(x)^{F}-\bm{k}^{F},\bm{\nu}(x)\rangle x-\langle x,\bm{\nu}(x)\rangle(\bm{\nu}^{F}(x)-\bm{k}^{F}).$
Here $\bm{\nu}^{F}-\bm{k}^{F}$ is actually the anisotropic normal vector field $\bm{\nu}^{\bar{F}}$ with respect to a new Minkowski norm $\bar{F}(\xi):=F(\xi)-\langle\xi,\bm{k}^{F}\rangle$. Thus $X(x)$ can be regarded as a vector triple product $\bm{\nu}\times(x\times\bm{\nu}^{\bar{F}})$ in the $3$-dimensional space spanned by $x,\bm{\nu}$ and $\bm{\nu}^{\bar{F}}$. By performing parallel translations along $\bm{\nu}^{F}-\bm{k}^{F}$, we further obtain the higher order Minkowski formulae.
Theorem 1.1.
Let $\mathbf{W}\subset\mathbb{R}^{n+1}$ be a classical wedge and $\Sigma\subset\overline{\mathbf{W}}$ be an immersed hypersurface. For $i=1,2$, we assume
$\langle\bm{\nu}^{F},\bm{n}_{i}\rangle=\omega_{0}^{i}\quad\text{on}\quad\partial_{i}\Sigma,$
where $\omega_{0}^{i}$ are constants. Suppose $\bm{k}^{F}$ is a constant vector in $\mathbb{R}^{n+1}$ satisfying
$\langle\bm{k}^{F},\bm{n}_{i}\rangle=\omega_{0}^{i},\quad\text{for}\quad i=1,2.$
Then we have, for $r\in\{1,\dots,n\},$
$\int_{\Sigma}H^{F}_{r-1}(F(\bm{\nu})-\langle\bm{k}^{F},\bm{\nu}\rangle)\,\mathrm{d}A=\int_{\Sigma}H_{r}^{F}\langle x,\bm{\nu}\rangle\,\mathrm{d}A.$ (1.1)
In particular, we have
$\int_{\Sigma}(F(\bm{\nu})-\langle\bm{k}^{F},\bm{\nu}\rangle)\,\mathrm{d}A=\int_{\Sigma}H^{F}\langle x,\bm{\nu}\rangle\,\mathrm{d}A.$ (1.2)
By carefully investigating the relationships of normal vectors on each part of $\partial\Sigma$, we first prove that the normal translation map $\zeta$ starting from the hypersurface $\Sigma$ can cover the domain enclosed by the hypersurface. Upon accomplishing this key step, similar to the closed hypersurface case, we can finally establish the following Heintze-Karcher inequality for anisotropic hypersurfaces with free boundary.
Theorem 1.2.
Let $\mathbf{W}\subset\mathbb{R}^{n+1}$ be a classical wedge and $\Sigma\subset\overline{\mathbf{W}}$ be a smooth, compact, embedded, strictly anisotropic mean convex $\bm{\omega}$-capillary hypersurface with $\omega^{i}(x)\leq 0$. Let $\Omega$ be the enclosed domain by $\Sigma$ and $\partial\mathbf{W}$. Then
$\int_{\Sigma}\frac{F(\bm{\nu})}{H^{F}}\mathrm{~{}d}A\geq(n+1)|\Omega|.$
Moreover, equality holds if and only if $\Sigma$ is an anisotropic free boundary truncated Wulff shape.
Remark 1.3.
Given a constant vector $\bm{k}_{0}$ defined in [12], we set $F(\xi):=|\xi|-\langle\bm{k}_{0},\xi\rangle$, then the above theorem reduces to Theorem 1.5 in [12] for $|\bm{k}_{0}|<1$.
Now we get the Alexandrov-type theorem for the free boundary case as follows.
Theorem 1.4.
Let $\Sigma\subset\overline{\mathbf{W}}$ be a smooth embedded compact anisotropic free boundary hypersurface with constant $r$-th anisotropic mean curvature for some $r\in\{1,...,n\}$. Then $\Sigma$ is an anisotropic free boundary truncated Wulff shape.
Inspired by the work in [26], we reduce the capillary case to the free boundary case via choosing a new Minkowski norm. We have the following analogous theorems for the capillary case.
Theorem 1.5.
For $i=1,2$, let $\omega_{0}^{i}\in(-F(-\bm{n}_{i}),F(\bm{n}_{i}))$ and $\mathbf{W}\subset\mathbb{R}^{n+1}$ be a classical wedge. Assume that a constant vector $\bm{k}^{F}$ satisfies
$\langle\bm{k}^{F},\bm{n}_{i}\rangle=\omega^{i}_{0}\,\text{for each}\,\bm{n}_{i}\quad\text{and}\quad F^{o}(\bm{k}^{F})<1.$
Let $\Sigma\subset\overline{\mathbf{W}}$ be a smooth, compact, embedded, strictly anisotropic mean convex hypersurface with $\langle\bm{\nu}^{F}(x),\bm{n}_{i}\rangle:=\omega^{i}(x)\leq\omega_{0}^{i}$ for $x\in\partial\Sigma\cap P_{i},\,i=1,2$. Let $\Omega$ be the enclosed domain by $\Sigma$ and $\partial\mathbf{W}$. Then
$\int_{\Sigma}\frac{F(\bm{\nu})-\left\langle\mathbf{\bm{\nu}},\bm{k}^{F}\right\rangle}{H^{F}}\mathrm{~{}d}A\geq(n+1)|\Omega|.$ (1.3)
Moreover, the equality holds if and only if $\Sigma$ is a $\bm{\omega}_{0}$-capillary truncated Wulff shape.
Theorem 1.6.
Let constants $\omega_{0}^{i}\in(-F(-\bm{n}_{i}),F(\bm{n}_{i}))$. Let $\Sigma\subset\overline{\mathbf{W}}$ be a smooth embedded compact anisotropic $\bm{\omega}_{0}$-capillary hypersurface with constant $r$-th anisotropic mean curvature for some $r\in\{1,...,n\}$. Then $\Sigma$ is an $\bm{\omega}_{0}$-capillary truncated Wulff shape.
Remark 1.7.
For convenience, we prove the free boundary case first and then derive the capillary case from the free boundary case. However, the technique of our proof can directly handle the capillary case, as shown in the subsequent proof.
The paper is organized as follows. In section 2, we will explain the anisotropic setting and geometric setting that will be used later; In section 3, we will prove the Minkowski-type formulae for the capillary boundary case in a wedge; In section 4, we prove the Heintze-Karcher inequality for the free boundary case; In section 5, we obtain the Alexandrov-type theorem for the free boundary case in a wedge, and then use a clever trick to directly derive the Alexandrov-type theorem for the general capillary case.
2. Preliminaries
2.1. Anisotropic setting
Let $F\colon\mathbb{S}^{n}\to\mathbb{R}$ is a $C^{2}$ positive function on $\mathbb{S}^{n}$, such that
$\left(\nabla^{2}F+F\sigma\right)>0,$
where $\sigma$ is the canonical metric on $\mathbb{S}^{n}$, $\nabla$ denote the gradient on $\mathbb{S}^{n}$ and $\nabla^{2}$ is the Hessian on $\mathbb{S}^{n}$. Let $A^{F}:=\nabla^{2}F+F\sigma$ such that
$\langle A^{F}\circ d\bm{\nu}(X),Y\rangle=\langle\nabla_{X}\bm{\nu}^{F},Y\rangle.$
The dual function $F^{o}$ of $F$ is defined by
$F^{o}(x)=\sup\left\{\frac{\langle x,\xi\rangle}{F(\xi)}\mid\xi\in\mathbb{S}^{n}\right\}.$
The Wulff shape centered at $x_{0}$ with radius $r_{0}$ is defined to be
$\mathcal{W}_{r_{0}}(x_{0})=\{x\in\mathbb{R}^{n+1}\mid F^{o}(x-x_{0})=r_{0}\}.$
$\mathcal{W}_{r_{0}}(x_{0})$ is a scaling of $\mathcal{W}_{1}(x_{0})$. It bounds a convex open set $D$. The Cahn-Hoffman map associated with $F$ is given by
$\Phi\colon\mathbb{S}^{n}\to\mathbb{R}^{n+1},\quad\Phi(\xi):=\nabla F(\xi)+F(\xi)\xi.$
Notice that the image $\Phi\left(\mathbb{S}^{n}\right)$ is a Wulff shape with radius $1$ centered at 00. Let $\tilde{F}$ be the positive one-homogeneous extension of $F$ to $\mathbb{R}^{n+1}$. We have $\Phi(x)=D\tilde{F}(x)$, where $D$ is the Euclidean derivative.
We collect some well-known facts on $F$ and $F^{o}$, see [10].
Proposition 2.1.
The following statements hold:
  1. (1)
    $F^{o}(\Phi(\xi))=1$, for $\xi\in S^{n}$.
  2. (2)
    $\left\langle\Phi(\xi),\xi\right\rangle=\tilde{F}(\xi)$ and $\left\langle DF^{o}(x),x\right\rangle=F^{o}(x)$.
  3. (3)
    $F^{o}(x)D_{\xi}\tilde{F}(DF^{o}(x))=x$ and $F(\xi)D_{x}F^{o}(D\tilde{F}(\xi))=\xi$.
  4. (4)
    $\langle\xi,y\rangle\leq\tilde{F}(\xi)F^{o}(y)$. The equality holds if and only if $y=\Phi(\xi)$.
2.2. Geometric setting
Suppose $\Sigma\subset\overline{\mathbf{W}}$ is a smooth compact embedded hypersurface, possibly containing corners. We require the corners to be of at most codimension $2$. In other words, $\Sigma$ is locally diffeomorphic to an open set in $\mathbb{R}_{\geq 0}^{2}\times\mathbb{R}^{n-2}$. The hypersurface $\Sigma$ can be decomposed as
$\Sigma=\Sigma_{0}\sqcup\Sigma_{1}\sqcup\Sigma_{2},$
where $\Sigma_{i}$ is a smooth manifold of dimension $n-i$. In this decomposition, $\Sigma_{0}$ is the interior of $\Sigma$ and $\Sigma_{1}\sqcup\Sigma_{2}$ is the boundary of $\Sigma$. We assume $\Sigma_{0}\subset\mathbf{W}$, $\Sigma_{1}\subset P_{1}\cup P_{2}$ and $\Sigma_{2}\subset L$.
In addition, the boundary $\partial\Sigma$ is a closed topological submanifold in $\partial\mathbf{W}$. Since $\partial\mathbf{W}$ is homeomorphic to $\mathbb{R}^{n}$, by Jordan’s theorem, $\partial\Sigma$ has a well-defined interior $\Gamma$ in $\partial\mathbf{W}$. $\Sigma\cup\Gamma$ is a closed topological submanifold in $\mathbb{R}^{n+1}$. Hence, again by Jordan’s theorem, there is a well-defined interior $\Omega$ of $\Sigma\cup\Gamma$ in $\mathbb{R}^{n+1}$.
Let $x$ be the position vector, $\bm{\nu}$ be the unit normal vector field on $\Sigma$ which points outside $\Omega$. For $i=1,2$, define $\partial_{i}\Sigma=\partial\Sigma\cap\overline{P_{i}}$. Then $\partial_{i}\Sigma$ is a smooth manifold with boundary $\Sigma_{0}$.
Regard $\partial_{i}\Sigma$ as a submanifold of $\Sigma$. This inclusion determines a unit normal vector field on $\partial_{i}\Sigma$ which points outside $\Sigma$. We denote it by $\bm{\mu}_{i}$.
Regard $\partial_{i}\Sigma$ as a submanifold of $P_{i}$. This inclusion determines a unit normal vector field on $\partial_{i}\Sigma$ which points inside $\Gamma$. We denote it by $\bm{m}_{i}$.
Regard $\Sigma_{2}$ as a submanifold of $\partial_{i}\Sigma$. This inclusion determines a unit normal vector field on $\Sigma_{2}$ which points outside $\partial_{i}\Sigma$. We denote it by $\bm{\tau}_{i}$.
Regard $\Sigma_{2}$ as a submanifold of $L$. This inclusion determines a unit normal vector field on $\Sigma_{2}$ which points inside $\Gamma\cap L$. We denote it by $\bm{l}$.
Fix $i=1$ or $2$. Then $\bm{\nu},\bm{n}_{i},\bm{\mu}_{i},\bm{m}_{i}$ all belong to a plane which is orthogonal to $\partial_{i}\Sigma$. $\bm{\nu},\bm{n}_{1},\bm{n}_{2},\bm{\tau}_{1},\bm{\tau}_{2},\bm{l}$ all belong to a $3-$dimensional linear subspace which is orthogonal to $\Sigma_{0}$.
We will always work in the above setting. It is illustrated in Figure 1.
Figure 1. An illustration of the vector fields defined above.
We have the following lemma of these vector fields.
Lemma 2.2.
If $\Sigma$ intersects with $P_{1},P_{2}$ and $L$ transversally, i.e.,$\bm{\nu},\bm{n}_{i}$ are linearly independent along $\partial_{i}\Sigma$ for $i=1,2$ and $\bm{\nu},\bm{n}_{1},\bm{n}_{2}$ are linearly independent along $L$, then
  1. (1)
    For $i=1,2$, $\bm{\mu}_{i}$ is characterized by the property that $\bm{\mu}_{i}$ is a unit vector orthogonal to $\partial_{i}\Sigma$ and
    $\langle\bm{\mu}_{i},\bm{\nu}\rangle=0,\quad\langle\bm{\mu}_{i},\bm{n}_{i}\rangle\geq 0.$
  2. (2)
    For $i=1,2$, $\bm{m}_{i}$ is characterized by the property that $\bm{m}_{i}$ is a unit vector orthogonal to $\partial_{i}\Sigma$ and
    $\langle\bm{m}_{i},\bm{n}_{i}\rangle=0,\quad\langle\bm{m}_{i},\bm{\nu}\rangle\leq 0.$
  3. (3)
    For $i=1,2$, $\bm{\tau}_{i}$ is characterized by the property that $\bm{\tau}_{i}$ is a unit vector orthogonal to $\Sigma_{0}$ and
    $\langle\bm{\tau}_{i},\bm{\nu}\rangle=\langle\bm{\tau}_{i},\bm{n}_{i}\rangle=0,\quad\langle\bm{\tau}_{i},\bm{n}_{3-i}\rangle\geq 0.$
  4. (4)
    $\bm{l}$ is characterized by the property that $\bm{l}$ is a unit vector orthogonal to $\Sigma_{0}$ and
    $\langle\bm{l},\bm{n}_{1}\rangle=\langle\bm{l},\bm{n}_{2}\rangle=0,\quad\langle\bm{l},\bm{\nu}\rangle\leq 0.$
Proof.
The proof is quite trivial. We will only prove (1) here. (2), (3) and (4) follow from almost the same arguments.
For (1), $\bm{\mu}_{i}$ is orthogonal to $\partial_{i}\Sigma$ and tangent to $\Sigma$ by definition. Since $\bm{\nu}$ is orthogonal to $\Sigma$, we have $\langle\bm{\mu}_{i},\bm{\nu}\rangle=0$. For each point $x\in\partial_{i}\Sigma$, the fact that $\bm{\mu}_{i}(x)$ points outside $\Sigma$ implies that there is a smooth curve $\alpha$ starting at $x$ with initial velocity $-\bm{\mu}_{i}(x)$ which lies entirely in $\Sigma$. Since $\Sigma\subset\overline{\mathbf{W}}$, $\alpha$ lies entirely in $\overline{\mathbf{W}}$. This implies that $\langle\bm{\mu}_{i},\bm{n}_{i}\rangle\geq 0$. It remains to show that there is exactly one vector field on $\partial_{i}\Sigma$ satisfying these properties. In fact, the transversality condition implies that $\bm{\nu},\bm{n}_{i}$ are linearly independent. The normal space of $\partial_{i}\Sigma$ is of dimension $2$. Hence, $\bm{\nu},\bm{n}_{i}$ form the basis of this normal space. As a result, $\bm{\mu}_{i}$ is a unit vector spanned by $\bm{\nu},\bm{n}_{i}$ and satisfies
$\langle\bm{\mu}_{i},\bm{\nu}\rangle=0,\quad\langle\bm{\mu}_{i},\bm{n}_{i}\rangle\geq 0.$
Such $\bm{\mu}_{i}$ is obviously unique. ∎
Remark 2.3.
If the transversality condition fails in the previous lemma, the properties still hold. However, these vector fields may not be uniquely determined by the above properties.
Proposition 2.4.
Let $\Sigma\subset\overline{\mathbf{W}}$ be a smooth, compact, embedded anisotropic $\bm{\omega}_{0}$-capillary hypersurface. Then $\omega_{0}^{i}\in(-F(-\bm{n}_{i}),\,F(\bm{n}_{i}))$ if and only if $\Sigma$ and the wedge $\mathbf{W}$ are transversal on $\Sigma_{1}$.
Proof.
By the proposition 2.1, we know
$-F\left(-\bm{n}_{i}\right)=-F\left(-\bm{n}_{i}\right)F^{o}(\Phi(\bm{\nu}))\leq\langle\Phi(\bm{\nu}),\bm{n}_{i}\rangle\leq F^{o}(\Phi(\bm{\nu}))F(\bm{n}_{i})=F(\bm{n}_{i}).$ (2.1)
The above inequality is strict, $\Sigma$ must intersect $\partial\mathbf{W}$ transversally. Otherwise, there exists a point $x$ such that $\bm{\nu}(x)=\pm\bm{n}_{i}$, then $\langle\Phi(\bm{\nu}),\bm{n}_{i}\rangle=-F(-\bm{n}_{i})$ or $F(\bm{n}_{i})$ at $x$. This contradicts the strict inequality.
If $\omega_{0}^{i}=-F(-\bm{n}_{i})$ or $F(\bm{n}_{i})$, then the equality sign of the inequality in (2.1) holds. By proposition 2.1 and the injectivity of the map $\Phi$, we obtain $\bm{\nu}=\pm\bm{n}_{i}$. This contradicts with transversality. ∎
Let $\omega_{0}^{1},\omega_{0}^{2}$ be constants and let $\bm{\omega}_{0}=(\omega_{0}^{1},\omega_{0}^{2})$. We call $\Sigma$ an anisotropic $\bm{\omega}_{0}$-capillary hypersurface in $\mathbf{W}$ if
$\langle\bm{\nu}^{F},\bm{n}\rangle=\omega_{0}^{i}\quad\text{on}\quad\partial_{i}\Sigma.$
Truncated Wulff shapes are always anisotropic capillary hypersurfaces in $\mathbf{W}$. In fact, suppose $\mathcal{W}_{r}(y)$ is a Wulff shape and intersects with $\partial_{i}\Sigma$ at some point $x$. The anisotropic normal of $\mathcal{W}_{r}(y)$ at $x$ is
$\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F}=\frac{x-y}{r}.$
Thus, $\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}_{i}\rangle=-\frac{1}{r}\langle y,-\bm{n}_{i}\rangle$ is constant. Conversely, for any constant $\bm{\omega}_{0}=(\omega_{0}^{1},\omega_{0}^{2})$, let $\bm{k^{F}}\in\mathbb{R}^{n+1}$ be a vector satisfying
$\langle\bm{k}^{F},\bm{n}_{i}\rangle=\omega_{0}^{i}\quad\text{for}\quad i=1,2.$
Consider the Wulff shape $\mathcal{W}_{1}(\bm{k}^{F})$. If $\mathcal{W}_{1}(\bm{k}^{F})\cap\partial_{i}\Sigma\neq\varnothing$ for $i=1,2$, then $\mathcal{W}_{1}(\bm{k}^{F})\cap\overline{\mathbf{W}}$ is an anisotropic $\bm{\omega}_{0}$-capillary hypersurface in $\mathbf{W}$.
3. Minkowski type formulae
In this section, we will first prove a Minkowski-type formula for anisotropic capillary hypersurface in a classical wedge then the higher order Minkowski-type formulae.
Proof of Theorem 1.1.
Consider the vector field on $\Sigma$ given by
$X(x)=\langle\bm{\nu}^{F}-\bm{k}^{F},\bm{\nu}\rangle x-\langle x,\bm{\nu}(x)\rangle(\bm{\nu}^{F}(x)-\bm{k}^{F}).$
Applying the divergence theorem, we obtain
$\int_{\Sigma}\operatorname{div}(X)=\int_{\partial_{1}\Sigma}\langle X,\bm{\mu}_{1}\rangle+\int_{\partial_{2}\Sigma}\langle X,\bm{\mu}_{2}\rangle.$ (3.1)
We claim that $\langle X,\bm{\mu}_{i}\rangle=0$ for $i=1,2$.
In fact, $X$ can be characterized by the property that
$\langle X,Y\rangle=\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge x,\bm{\nu}\wedge Y\rangle,\quad\text{for all}\quad Y\in T\Sigma.$
Then
$\langle X,\bm{\mu}_{i}\rangle=\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge x,\bm{\nu}\wedge\bm{\mu}_{i}\rangle.$
Along $\partial_{i}\Sigma$, $\bm{\nu}^{F}-\bm{k}^{F}$ and $x$ are always orthogonal to $\bm{n}_{i}$, while $\bm{n}_{i}$ can be spanned by $\bm{\nu}$ and $\bm{\mu}_{i}$. This implies that
$\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge x,\bm{\nu}\wedge\bm{\mu}_{i}\rangle=0.$
Thus (3.1) reduces to
$\int_{\Sigma}\operatorname{div}(X)=0.$
Next, we will show that $\operatorname{div}(X)=n(F(\bm{\nu})-\langle\bm{\nu},\bm{k}^{F}\rangle-H^{F}\langle\bm{\nu},x\rangle)$.
Suppose $e_{1},\ldots,e_{n}$ is any local frame on $\Sigma$. Then
$\operatorname{div}(X)=g^{ij}\langle\nabla_{e_{i}}X,e_{j}\rangle.$
With the above characterization of $X$, we have
$\langle\nabla_{e_{i}}X,e_{j}\rangle$ $=\langle D_{e_{i}}X,e_{j}\rangle$
$=e_{i}\langle X,e_{j}\rangle-\langle X,D_{e_{i}}e_{j}\rangle$
$=\langle(D_{e_{i}}\bm{\nu}^{F})\wedge x,\bm{\nu}\wedge e_{j}\rangle+\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge e_{i},\bm{\nu}\wedge e_{j}\rangle$
$+\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge x,(D_{e_{i}}\bm{\nu})\wedge e_{j}\rangle.$
The third equality uses the fact that $D_{e_{i}}\bm{k}^{F}=0$ and $D_{e_{i}}x=e_{i}$. We compute these three terms separately.
$g^{ij}\langle(D_{e_{i}}\bm{\nu}^{F})\wedge x,\bm{\nu}\wedge e_{j}\rangle=-g^{ij}\langle D_{e_{i}}\bm{\nu}^{F},e_{j}\rangle\langle\bm{\nu},x\rangle=-nH^{F}\langle\bm{\nu},x\rangle.$
The first equality follows from $\langle D_{e_{i}}\bm{\nu}^{F},\bm{\nu}\rangle=0$.
$g^{ij}\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge e_{i},\bm{\nu}\wedge e_{j}\rangle=g^{ij}\langle\bm{\nu}^{F}-\bm{k}^{F},\bm{\nu}\rangle\langle e_{i},e_{j}\rangle=n\langle\bm{\nu}^{F}-\bm{k}^{F},\bm{\nu}\rangle=n(F(\bm{\nu})-\langle\bm{k}^{F},\bm{\nu}\rangle).$
The first equality follows from $\langle e_{i},\bm{\nu}\rangle=0$.
$g^{ij}\,(D_{e_{i}}\bm{\nu})\wedge e_{j}=g^{ij}h_{i}^{k}\,e_{k}\wedge e_{j}=h^{jk}\,e_{k}\wedge e_{j}=0.$
The last equality follows from the fact that $h$ is symmetric. Thus,
$g^{ij}\langle(\bm{\nu}^{F}-\bm{k}^{F})\wedge x,(D_{e_{i}}\bm{\nu})\wedge e_{j}\rangle=0.$
From the above computations,
$\operatorname{div}(X)=n(F(\bm{\nu})-\langle\bm{\nu},\bm{k}^{F}\rangle-H^{F}\langle\bm{\nu},x\rangle).$
This completes the proof of (1.1) for $r=1$.
Next, for small $t$, we define
$\psi(x,t)=x+t(\bm{\nu}^{F}(x)-\bm{k}^{F})\quad\text{for}\quad x\in\Sigma.$
which is a family of parallel hypersurfaces $\Sigma_{t}$
On one hand, the $\bm{\omega}$-capillarity condition and the definition of $\bm{k}^{F}$ yield that for any $x\in\partial\Sigma\cap P_{i}$,
$\langle x+t(\bm{\nu}^{F}(x)-\bm{k}^{F}),\bm{n}_{i}\rangle=t(\omega^{i}-\omega^{i})=0.$
Hence, $\psi_{t}(x)\in\partial\mathbf{W}$ for $x\in\partial\Sigma$ which means $\partial\Sigma_{t}\subset\partial\mathbf{W}$.
On the other hand, if $e_{1}^{F},\cdots,e_{n}^{F}$ are anisotropic principal directions at $x\in\Sigma$ corresponding to $\kappa_{i}^{F}$ for $i=1,\ldots,n$, we have
$(\psi_{t})_{*}(e_{i}^{F})=(1+t\kappa_{i}^{F})e_{i}^{F},\quad i=1,\ldots,n.$
These imply $\bm{\nu}_{\Sigma_{t}}(\psi_{t}(x))=\bm{\nu}(x)$, so $\bm{\nu}^{F}_{\Sigma_{t}}(\psi_{t}(x))=\bm{\nu}^{F}(x)$. Here $\bm{\nu}_{\Sigma_{t}}$ and $\bm{\nu}^{F}_{\Sigma_{t}}$ denote the outward normal and anisotropic normal to $\Sigma_{t}$ respectively. Moreover, we have
$\langle\bm{\nu}^{F}_{\Sigma_{t}}(\psi_{t}(x)),\bm{n}_{i}\rangle=\langle\bm{\nu}^{F}(x),\bm{n}_{i}\rangle=\omega^{i}.$
Therefore, $\Sigma_{t}$ is also an anisotropic $\bm{\omega}$-capillary hypersurface in $\overline{\mathbf{W}}$ for any small $t$. So using (1.2) for every such $\Sigma_{t}$, we find that
$\int_{\Sigma_{t}=\psi_{t}(\Sigma)}(F(\bm{\nu}_{t})-\langle\bm{k}^{F},\bm{\nu}_{t}\rangle)-H^{F}(t)\langle\psi_{t},\bm{\nu}_{t}\rangle\mathrm{d}A_{t}=0.$ (3.2)
It is easy to see that the corresponding anisotropic principal curvatures are given by
$\kappa_{i}^{F}(\psi_{t}(x))=\frac{\kappa_{i}^{F}(x)}{1+t\kappa_{i}^{F}(x)}.$
Hence, if we denote $\mathcal{P}_{n}(t)$ by
$\mathcal{P}_{n}(t)=\prod_{i=1}^{n}(1+t\kappa_{i}^{F})=\sum_{i=0}^{n}\binom{n}{i}H_{i}^{F}t^{i},$
then the anisotropic mean curvature of $\Sigma_{t}$ at $\psi_{t}(x)$ is given by
$H^{F}(t)=\frac{\mathcal{P}_{n}^{\prime}(t)}{\mathcal{P}_{n}(t)}=\frac{\sum\limits_{i=0}^{n}i\displaystyle\binom{n}{i}H_{i}^{F}t^{i-1}}{\mathcal{P}_{n}(t)}.$
Therefore, combining the tangential Jacobian of $\psi_{t}$ along $\Sigma$ at $x$
$\mathcal{J}^{\Sigma}\psi_{t}(x)=\prod_{i=1}^{n}(1+t\kappa_{i}^{F}(x))=\mathcal{P}_{n}(t).$
We can apply (3.2) to get
$\int_{\Sigma}(F(\bm{\nu})-\langle\bm{k}^{F},\bm{\nu}\rangle)\mathcal{P}_{n}(t)-\mathcal{P}_{n}^{\prime}(t)(\langle x,\bm{\nu}\rangle+t\langle\bm{\nu}^{F},\bm{\nu}\rangle-t\langle\bm{k}^{F},\bm{\nu}\rangle)\,\mathrm{d}A_{x}=0.$
Hence, by a direct computation, we obtain (1.1). ∎
Remark 3.1.
The same proof is also valid for general wedges bounded by multiple hyperplanes; for a detailed definition, refer to [12].
4. Heintze-Karcher type inequality
In this section, to establish Theorem 1.2, we first prove the following monotonicity lemma.
Lemma 4.1.
Let $z$ be a point in $S^{n}$ and $\gamma\colon[0,\pi]\to S^{n}$ be a unit speed geodesic with $\gamma(0)=z$. Then
$f(t)=\langle\Phi(\gamma(t)),z\rangle$
is a strictly decreasing function on $[0,\pi]$.
Proof.
We embed $S^{n}$ in $\mathbb{R}^{n+1}$ and let $z$ also denote the position vector of itself. Then
$f^{\prime}(t)=\langle D_{\gamma^{\prime}(t)}\Phi(\gamma(t)),z\rangle=\langle A^{F}_{\gamma(t)}\gamma^{\prime}(t),z\rangle=\langle A^{F}_{\gamma(t)}\gamma^{\prime}(t),z^{\perp}\rangle,$ (4.1)
where $z^{\perp}$ is the orthogonal projection of $z$ onto $T_{\gamma(t)}S^{n}$. The last equality follows from the fact that $A^{F}_{\gamma(t)}$ is an endomorphism of $T_{\gamma(t)}S^{n}$. Since $z$ always lies in the plane spanned by $\gamma(t)$ and $\gamma^{\prime}(t)$, we have
$z=\langle z,\gamma(t)\rangle\gamma(t)+\langle z,\gamma^{\prime}(t)\rangle\gamma^{\prime}(t).$
Hence, $z^{\perp}=\langle z,\gamma^{\prime}(t)\rangle\gamma^{\prime}(t)$. Go back to (4.1),
$f^{\prime}(t)=\langle z,\gamma^{\prime}(t)\rangle\langle A^{F}_{\gamma(t)}\gamma^{\prime}(t),\gamma^{\prime}(t)\rangle=\langle z,\gamma^{\prime}(t)\rangle(\nabla^{2}F+F\sigma)(\gamma^{\prime}(t),\gamma^{\prime}(t)),$
where $\nabla^{2}F+F\sigma$ is positive. Notice $\gamma(t)$ can be written as
$\gamma(t)=\cos t\,z+\sin t\,\gamma^{\prime}(0).$
Thus,
$\gamma^{\prime}(t)=-\sin t\,z+\cos t\,\gamma^{\prime}(0).$
Then
$\langle z,\gamma^{\prime}(t)\rangle=-\sin t.$
It means that $f^{\prime}(t)<0$ on $(0,\pi)$, which proves the lemma. ∎
Proof of Theorem 1.2.
For any $x\in\Sigma$, let $\kappa_{i}^{F}(x)$ be the anisotropic principal curvatures and $e_{i}^{F}(x)$ be the anisotropic principal directions. Since $H^{F}>0$,
$\max_{i}\kappa_{i}^{F}(x)\geq H^{F}(x)>0,\quad\text{for all}\quad x\in\Sigma.$
Define
$Z=\left\{(x,t)\in\Sigma\times\mathbb{R}\mid 0<t\leq\frac{1}{\max\limits_{i}\kappa_{i}^{F}(x)}\right\}$
and
$\zeta\colon Z\to\mathbb{R}^{n+1},\quad(x,t)\mapsto x-t\bm{\nu}^{F}(x).$
We claim that $\Omega\subset\zeta(Z)$. For any $y\in\Omega$, we consider a family of Wulff shapes $\{\mathcal{W}_{r}(y)\}_{r>0}$ which forms a foliation of $\mathbb{R}^{n+1}\setminus\{y\}$. Let
$r_{0}=\inf\{r>0\mid\mathcal{W}_{r}(y)\cap\Sigma\neq\varnothing\}.$
Since $\mathcal{W}_{r}(y)\subset\Omega$ when $r$ is small, it follows that $r_{0}>0$. Intuitively, this can be seen as a process where the Wulff shape continues to grow larger and larger until it comes into contact with $\Sigma$. And $r_{0}$ represents the time when $\mathcal{W}_{r}(y)$ first touches $\Sigma$. Let $D$ be the open region enclosed by $\mathcal{W}_{r_{0}}(y)$. We have
$D\cap\mathbf{W}\subset\Omega.$ (4.2)
This inclusion is crucial in later discussions.
Let $x$ be a point in $\mathcal{W}_{r_{0}}(y)\cap\Sigma$ and $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ be the outwrad unit normal vector of $\mathcal{W}_{r_{0}}(y)$ at $x$. We have
$\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F}=\frac{x-y}{r_{0}}.$ (4.3)
Next, we will always work in the tangent space $T_{x}\mathbb{R}^{n+1}$. For simplicity, we will write the vector $\bm{\nu}(x)$ as $\bm{\nu}$. The same simplification applies to $\bm{n}_{i}(x),\bm{\mu}_{i}(x),\bm{m}_{i}(x),\bm{\tau}_{i}(x)$ or $\bm{l}(x)$, if they exist. There are three cases.
Case 1. $x\in\Sigma_{0}$. In this case, $\mathcal{W}_{r_{0}}(y)$ is tangent to $\Sigma_{0}$, i.e., $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}=\bm{\nu}$. Therefore, $\bm{\nu}^{F}=\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F}=\frac{x-y}{r_{0}}$. By the inclusion (4.2), we have
$\kappa_{i}^{F}\leq\frac{1}{r_{0}},\quad\text{for all}\quad i=1,\ldots,n.$
Hence, $(x,r_{0})\in Z$ and $y=\zeta(x,r_{0})\in\zeta(Z)$.
Case 2. $x\in\Sigma_{1}$. Suppose $x\in\partial_{i}\Sigma$. Then $\mathcal{W}_{r_{0}}(y)$ is tangent to $\partial_{i}\Sigma$. This implies that $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the plane spanned by $\bm{\nu}$ and $\bm{\mu}_{i}$. Hence, $\bm{\nu},\bm{n}_{i},\bm{\mu}_{i},\bm{m}_{i},\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ all belong to a same plane. On one hand, by (4.3),
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}_{i}\rangle=-\frac{\langle y,\bm{n}_{i}\rangle}{r_{0}}>0.$ (4.4)
On the other hand, by the inclusion (4.2), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{\mu}_{i}\rangle\leq 0\quad\text{and}\quad\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{m}_{i}\rangle\leq 0.$ (4.5)
[…]
Figure 2. $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ must lies on the red arc.
Therefore, by set $z=\bm{n}_{i}$ in Lemma 4.1, we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}_{i}\rangle\leq\langle\bm{\nu}^{F},\bm{n}_{i}\rangle=\omega^{i}\leq 0,$
which contradicts with (4.4). This means that case 2 can not happen.
Case 3. $x\in\Sigma_{2}$. In this case, $\mathcal{W}_{r_{0}}(y)$ is tangent to $\Sigma_{2}$. Hence, $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the normal space of $\Sigma_{2}$, which is a $3-$dimensional linear subspace containing $\bm{\nu},\bm{n}_{1},\bm{n}_{2},\bm{\tau}_{1},\bm{\tau}_{2},\bm{l}$.
On one hand, by (4.3), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}_{i}\rangle>0,\quad\text{for}\quad i=1,2.$ (4.6)
On the other hand, by the inclusion (4.2), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{\tau}_{i}\rangle\leq 0\quad\text{for}\quad i=1,2\quad\text{and}\quad\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{l}\rangle\leq 0.$ (4.7)
This follows from the same argument as in case 2. If we regard $\bm{\nu},\bm{n}_{1},\bm{n}_{2},\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ as points in $S^{2}$, (4.7) is equivalent to the statement that $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the geodesic triangle with vertices $\bm{\nu},-\bm{n}_{1},-\bm{n}_{2}$ whose three sides are all length minimizing. See Figure 3 for an illustration.
Figure 3. $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ must lies on the geodesic triangle with vertices $-\bm{n}_{1},-\bm{n}_{2}$ and $\bm{\nu}$. The vector $\bm{l}$ does not occur in this figure because it points perpendicularly into the paper.
Denote this geodesic triangle by $\Delta$. We also denote the length minimizing geodesic connecting $\bm{n}_{1}$ and $\bm{n}_{2}$ by $\gamma$ and suppose $\eta$ is a geodesic connecting $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ and $\bm{\nu}$ which lies inside $\Delta$. If we extend $\eta$, it will intersect with $\gamma$ at a certain point $\bm{n}$. Moreover, $\bm{n}$ can be written as
$\bm{n}=\lambda_{1}\bm{n}_{1}+\lambda_{2}\bm{n}_{2}\quad\text{with}\quad\lambda_{1},\lambda_{2}\geq 0.$
By (4.6), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}\rangle>0.$
However, if we apply Lemma 4.1 with $z=\bm{n}$ to $\eta$, we obtain
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}\rangle\leq\langle\bm{\nu}^{F},\bm{n}\rangle=\lambda_{1}\omega_{1}+\lambda_{2}\omega_{2}\leq 0.$
This is a contradiction. Therefore, case 3 can not happen either. The claim is thus proved.
Next, by direct computations, the Jacobian of $\zeta$ along $Z$ at $(x,t)$ is
$J\zeta(x,t)=F(\bm{\nu})\prod_{i=1}^{n}(1-t\kappa_{i}^{F}).$
By $\Omega\subset\zeta(Z)$, the area formula yields
$|\Omega|\leq\int_{Z}J\zeta\,\mathrm{d}t\,\mathrm{d}A$ $=\int_{\Sigma}\int_{0}^{\frac{1}{\max\limits_{i}{\kappa_{i}^{F}}}}F(\bm{\nu})\prod_{i=1}^{n}(1-t\kappa_{i}^{F})\,\mathrm{d}t\,\mathrm{d}A$
$\leq\int_{\Sigma}\int_{0}^{\frac{1}{\max\limits_{i}{\kappa_{i}^{F}}}}F(\bm{\nu})(1-tH^{F})^{n}\,\mathrm{d}t\,\mathrm{d}A$
$\leq\int_{\Sigma}\int_{0}^{\frac{1}{H}}F(\bm{\nu})(1-tH^{F})^{n}\,\mathrm{d}t\,\mathrm{d}A$
$=\int_{\Sigma}\frac{F(\bm{\nu})}{(n+1)H^{F}}\,\mathrm{d}A.$
The second inequality comes from the AM-GM inequality. The third inequality follows from the fact that $\max\limits_{i}{\kappa_{i}^{F}}\geq H^{F}>0$. If the equality holds, then $\kappa_{1}^{F}(x)=\cdots=\kappa_{n}^{F}(x)$ for all $x\in\Sigma$. It follows from [9, 22] that $\Sigma$ must be a part of a Wulff shape $\mathcal{W}_{r}(y)$. If $\partial_{i}\Sigma\neq\varnothing$, suppose $x$ is a point in $\partial_{i}\Sigma$. Then
$\langle\bm{\nu}^{F}(x),\bm{n}_{i}\rangle=\langle\frac{x-y}{r},\bm{n}_{1}\rangle=-\frac{\langle y,\bm{n}_{i}\rangle}{r},$
which is a constant. Denote it by $\omega^{i}$. We have assumed that $\omega^{i}\leq 0$. If $\omega^{i}<0$, it is easy to check that the difference of $\Omega$ and $\zeta(Z)$ have a positive measure, which violates the equality in Heintze-Karcher inequality. Hence, $\omega^{i}=0$ i.e., $\Sigma$ is a truncated Wulff shape with a free boundary. ∎
Remark 4.2.
The theorem relies on the existence of a constant vector $\bm{k}^{F}$ with
$\langle\bm{k}^{F},\bm{n}_{i}\rangle=\omega_{i}\ \text{for}\ i=1,2\quad\text{and}\quad F^{o}(\bm{k}^{F})<1.$
Such $\bm{k}^{F}$ does not exist in general. In fact, we have the following proposition.
Proposition 4.3.
The constant vector $\bm{k}^{F}$ required by Theorem 1.5 exists if and only if the $\bm{\omega}_{0}$-capillary truncated Wulff shape in $\mathbf{W}$ intersects with $L$ transversally.
Proof.
Suppose such $\bm{k}^{F}$ exists. Consider the Wulff shape $\mathcal{W}_{1}(-\bm{k}^{F})$. Since $F^{o}(\bm{k}^{F})<1$, $O$ lies in the interior of $\mathcal{W}_{1}(\bm{k}^{F})$. This implies that $\mathcal{W}_{1}(\bm{k}^{F})$ intersects with $L$. Direct computation shows that $\mathcal{W}_{1}(-\bm{k}^{F})$ is an $\bm{\omega}_{0}$-capillary truncated Wulff shape. The intersection must be transversal. Otherwise, $L$ is tangent to $\mathcal{W}_{1}(-\bm{k}^{F})$. By the convexity of $\mathcal{W}_{1}(-\bm{k}^{F})$, $L$ cannot touch the interior of $\mathcal{W}_{1}(-\bm{k}^{F})$.
Suppose there is an $\bm{\omega}_{0}$-capillary truncated Wulff shape in $\mathbf{W}$ which intersects with $L$ transversally. Denote this Wulff shape by $\mathcal{W}_{0}$. Let $x_{0}$ be a point of $\mathcal{W}_{0}\cap L$, $\bm{\nu}_{0}$ be the outward unit normal vector of $\mathcal{W}_{0}$ at $x_{0}$ and $\bm{\nu}_{0}^{F}=\Phi(\bm{\nu}_{0})$. Then
$\langle\bm{\nu}_{0}^{F},\bm{n}_{i}\rangle=\omega^{i}_{0}\ \text{for}\ i=1,2\quad\text{and}\quad F^{o}(\bm{\nu}_{0}^{F})=1.$
By transversality, there is a nonzero vector $\bm{n}$ with
$\langle\bm{n},\bm{n}_{i}\rangle=0\ \text{for}\ i=1,2\quad\text{and}\quad\langle\bm{n},\bm{\nu}_{0}\rangle\neq 0.$
We have
$\frac{\partial F^{o}(\bm{\nu}_{0}^{F}+t\bm{n})}{\partial t}\bigg{|}_{t=0}=\langle\bm{n},\nabla F^{o}(\bm{\nu}_{0}^{F})\rangle=\frac{1}{F(\bm{\nu}_{0})}\langle\bm{n},\bm{\nu}_{0}\rangle\neq 0.$
Hence, we can find a small $t$ such that
$F^{o}(\bm{\nu}_{0}^{F}+t\bm{n})<F^{o}(\bm{\nu}_{0}^{F})=1.$
Then $\bm{\nu}_{0}^{F}+t\bm{n}$ is the desired $\bm{k}^{F}$. ∎
5. Alexandrov type theorem
In this section, we will start by using the results in section 4 to prove the Alexandrov theorem for the free boundary case, and then introduce a very useful idea and combine it with the above results to directly obtain the corresponding result in the case of capillary hypersurfaces.
Firstly, we prove that for the CMC hypersurface, the high-order mean curvature must be positive. The following proof also applies to the Capillary case.
Proposition 5.1.
Let $\omega^{i}\in(-F(-\bm{n}_{i}),F(\bm{n}_{i}))$ and $\Sigma\subset\overline{\mathbf{W}}$ be a smooth compact embedded anisotropic free boundary hypersurface, then $\Sigma$ has at least one point, at which all the anisotropic principal curvatures are positive.
Proof.
We fix a point $y\in L$. Consider the family of Wulff shapes $\mathcal{W}_{r}(y)$ centred at $y$. It is important to note that for any point $x\in\partial\Sigma\cap\partial\mathcal{W}_{r}(y)\subset\partial\mathbf{W}$, the following relationship holds:
$\langle\bm{\nu}^{F}_{\scriptscriptstyle\mathcal{W}}(x),\bm{n}_{i}\rangle=\left\langle\frac{x-y}{r},\bm{n}_{i}\right\rangle=0=\langle\bm{\nu}^{F}(x),\bm{n}_{i}\rangle.$ (5.1)
From $\Sigma$ is compact, there exists a sufficiently large radius $r$ such that $\Sigma$ is entirely enclosed by the Wulff shape $\mathcal{W}_{r}(y)$. Consequently, we can find the smallest radius $r_{0}>0$ at which $\mathcal{W}_{r_{0}}(y)$ makes initial contact with $\Sigma$ from the exterior at a point $x_{0}\in\Sigma$. Let $D$ be the open region enclosed by $\mathcal{W}_{r_{0}}(y)$. We have
$\Omega\subset D\cap\mathbf{W}.$ (5.2)
Now, there are three cases regarding the position of $x_{0}$. The proof is similar to the analysis in the Heintze-Karcher inequality above, and we will briefly describe it here.
Case 1. $x_{0}\in\Sigma_{0}$. Then $\Sigma$ and $\mathcal{W}_{r}(y)$ are tangent at $x_{0}$.
Case 2. $x_{0}\in\Sigma_{1}$. Suppose $x_{0}\in\partial_{i}\Sigma$. Then $\mathcal{W}_{r_{0}}(y)$ is tangent to $\partial_{i}\Sigma$. This implies that $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the plane spanned by $\bm{\nu}$ and $\bm{\mu}_{i}$. Hence, $\bm{\nu},\bm{n}_{i},\bm{\mu}_{i},\bm{m}_{i},\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ all belong to a same plane.
By the inclusion (5.2), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{\mu}_{i}\rangle\geq 0\quad\text{and}\quad\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{m}_{i}\rangle\leq 0.$ (5.3)
[…]
Figure 4. $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ must lies on the red arc.
Therefore, by set $z=\bm{n}_{i}$ in Lemma 4.1 and (5.1), we have
$\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F}(x_{0})=\bm{\nu}^{F}(x_{0}).$
This means that $\Sigma$ and $\mathcal{W}_{r}\left(y\right)$ are tangent at $x_{0}$.
Case 3. $x_{0}\in\Sigma_{2}$. In this case, $\mathcal{W}_{r_{0}}(y)$ is tangent to $\Sigma_{2}$. Hence, $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the normal space of $\Sigma_{2}$, which is a $3-$dimensional linear subspace containing $\bm{\nu},\bm{n}_{1},\bm{n}_{2},\bm{\tau}_{1},\bm{\tau}_{2},\bm{l}$.
By the inclusion (5.2), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{\tau}_{i}\rangle\geq 0\quad\text{and}\quad\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}},\bm{l}\rangle\leq 0.$ (5.4)
If we regard $\bm{\nu},\bm{n}_{1},\bm{n}_{2},\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ as points in $S^{2}$, (5.4) is equivalent to the statement that $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ lies in the geodesic triangle with vertices $\bm{\nu},\bm{n}_{1},\bm{n}_{2}$ whose three sides are all length minimizing.
Denote this geodesic triangle by $\Delta$. Also denote the length minimizing geodesic connecting $\bm{n}_{1}$ and $\bm{n}_{2}$ by $\gamma$. Suppose $\eta$ is a geodesic connecting $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ and $\bm{\nu}$ which lies inside $\Delta$. If we extend $\eta$, it will intersect with $\gamma$ at some point $\bm{n}$. See Figure 5 for an illustration.
Figure 5. $\bm{\nu}_{\scriptscriptstyle\mathcal{W}}$ must lies on the geodesic triangle with vertices $\bm{n}_{1},\bm{n}_{2}$ and $\bm{\nu}$.
$\bm{n}$ can be written as
$\bm{n}=\lambda_{1}\bm{n}_{1}+\lambda_{2}\bm{n}_{2}\quad\text{with}\quad\lambda_{1},\lambda_{2}\geq 0.$
By (5.1), we have
$\langle\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F},\bm{n}\rangle=0.$
However, if we apply Lemma 4.1 with $z=\bm{n}$ to $\eta$, we obtain
$\bm{\nu}_{\scriptscriptstyle\mathcal{W}}^{F}=\bm{\nu}^{F}.$
Therefore, $\Sigma$ and $\mathcal{W}_{r}\left(y\right)$ are tangent at $x_{0}$.
In both cases, by a similar argument as in the proof of Theorem 1.2, we have that the anisotropic principal curvatures of $\Sigma$ at $x_{0}$ are larger than or equal to $\frac{1}{r_{0}}$. ∎
Proof of Theorem 1.4.
Following the technique of Montiel and Ros in [20], by Proposition 5.1, we know that $H_{j}^{F}$ are positive, for $j\leq r$ and for any $x\in\Sigma$. Applying Theorem 1.2 and using the Maclaurin inequality $H_{1}^{F}\geq\left(H_{r}^{F}\right)^{1/r}$ and $H_{r}^{F}=constant$, we have
$(n+1)\left(H_{r}^{F}\right)^{1/r}|\Omega|\leq\left(H_{r}^{F}\right)^{1/r}\int_{\Sigma}\frac{F(\nu)}{H_{1}^{F}}\mathrm{~{}d}A\leq\int_{\Sigma}F(\nu)\mathrm{d}A.$
On the other hand, using Theorem 1.1 and the Maclaurin inequality $H_{r-1}^{F}\geq\left(H_{r}^{F}\right)^{\frac{r-1}{r}}$, we have
00 0 $=\int_{\Sigma}H_{r-1}^{F}F(\nu)-H_{r}^{F}\langle x,\nu\rangle\mathrm{d}A\geq\int_{\Sigma}\left(H_{r}^{F}\right)^{\frac{r-1}{r}}F(\nu)-H_{r}^{F}\langle x,\nu\rangle\mathrm{d}A$
$=\left(H_{r}^{F}\right)^{\frac{r-1}{r}}\int_{\Sigma}F(\nu)-\left(H_{r}^{F}\right)^{\frac{1}{r}}\langle x,\nu\rangle\mathrm{d}A=\left(H_{r}^{F}\right)^{\frac{r-1}{r}}\int_{\Sigma}F(\nu)\mathrm{d}A-(n+1)H_{r}^{F}|\Omega|.$
Thus equality in Theorem 1.2 holds, and hence $\Sigma$ is an anisotropic $\omega_{0}$-capillary Wulff shape. This completes the proof. ∎
Proof of Theorems 1.5 and 1.6.
Consider
$\bar{F}(\xi):=F(\xi)-\langle\xi,\bm{k}^{F}\rangle.$
Then
$\bm{\nu}^{\bar{F}}=\bar{F}(\bm{\nu})\bm{\nu}+\nabla^{\mathbb{S}^{n}}\bar{F}(\bm{\nu})=(F(\bm{\nu})-\langle\bm{\nu},\bm{k}^{F}\rangle)\bm{\nu}+\nabla^{\mathbb{S}^{n}}F(\bm{\nu})-(\bm{k}^{F})^{T}=\bm{\nu}^{F}-\bm{k}^{F}.$
Moreover, (i) $\bar{F}$ is a Minkowski norm.
In fact, it is easy to check that $\bar{F}\in C^{\infty}(\mathbb{R}^{n+1}\backslash\{0\})$, is a 1-homogeneous convex function satisfying $\bar{F}(\xi)>0$ (due to $F^{o}(\bm{k}^{F})<1$) when $x\neq 0$ and $\bar{F}(0)=0$.
Furthermore,
$A_{\bar{F}}(x):=\nabla^{\mathbb{S}^{n}}\nabla^{\mathbb{S}^{n}}\bar{F}(x)+\bar{F}(x)\sigma>0\text{ for }x\in\mathbb{S}^{n},$
the component is
$\bar{F}_{ij}+\bar{F}\sigma_{ij}=F_{ij}+F\sigma_{ij}.$
So $A_{\bar{F}}$ is positive definite matrix in $\mathbb{S}^{n}.$
(ii) $H_{\bar{F}}=\mathrm{div}\bm{\nu}_{\bar{F}}=\mathrm{div}\bm{\nu}^{F}=H_{F}.$
Hence, $\langle\bm{\nu}^{F},\bm{n}_{i}\rangle=\omega^{i}\text{ on }\partial\Sigma$ and $\langle\bm{k}^{F},\bm{n}_{i}\rangle=\omega_{0}^{i}$ implies that the free boundary condition
$\langle\bm{\nu}^{\bar{F}},\bm{n}_{i}\rangle=\langle\bm{\nu}^{F},\bm{n}_{i}\rangle-\langle k^{F},\bm{n}_{i}\rangle\leq 0.$
This means that $\Sigma$ is an embedded compact strictly anisotropic mean convex anisotropic $\bar{\bm{\omega}}$-capillary hypersurface with $\bar{\omega}^{i}(x)\leq 0$ in the wedge, and the anisotropic norm is $\bar{F}$.
From the inequality in Theorem 1.2, we immediately obtain the inequality in Theorem 1.5. As for the equalities in Theorem 1.5 and Theorem 1.6, we have known that $\Sigma$ is an anisotropic free boundary truncated Wulff shape associated with $\bar{F}$.
Moreover, from
$\bar{\Phi}(\mathbb{S}^{n})+\bm{k}^{F}=\Phi(\mathbb{S}^{n}),$
we obtain $\Sigma$ is an anisotropic $\bm{\omega}_{0}$-capillary truncated Wulff shape associated with $F$. ∎
Remark 5.2.
The above technique can be used to prove the Heintze-Karcher type inequality and Alexandrov-type theorem for anisotropic capillary hypersurfaces in some types of circular cones.
Acknowledgments.
The authors would like to thank Zhenxiao Xie and Chao Qian for their helpful comments. The authors were partially supported by the National Natural Science Foundation of China under grants No. 11831005 and No. 12061131014.